OpenStudy (anonymous):

PLEASE HELP ME!! REALLY NEED HELP!!! PLEASE!! For his exercise today, Eric plans to both run and swim. Let r be the number of laps he runs and let s be the number of laps he swims.Each lap he runs takes him 4 minutes, and each lap he swims takes him 3 minutes. He wants to exercise for more than 30 minutes today. Using the values and variables given, write an inequality describing this.

TheSmartOne (thesmartone):

\[4r+3s \ge30\]

OpenStudy (librarian):

^ Correct

TheSmartOne (thesmartone):

Oh \[4r+3s >30\] Wrote the wrong symbol...

TheSmartOne (thesmartone):

@babyblue123

TheSmartOne (thesmartone):

Ok so to answer your question...

TheSmartOne (thesmartone):

\[9+3x \ge t\]

TheSmartOne (thesmartone):

Ugh dang it...

TheSmartOne (thesmartone):

So sorry about it... The next one will be correct..

TheSmartOne (thesmartone):

http://www2.ccsd.ws/sbfaculty/team8e/jecole/Math/Inequalities%20symbols.htm Going to use it to make sure the symbol is correct

TheSmartOne (thesmartone):

\[9+3x \le t\]

TheSmartOne (thesmartone):

\[2x+3y >90\]

TheSmartOne (thesmartone):

\[3x >5y\]

TheSmartOne (thesmartone):

\[13t+11(106-t)=total~earned\]

TheSmartOne (thesmartone):

Degree measure of fourth angle = \[360-(90+5y+3+3y-9)\] or \[360-90-(5y+3)-(3y-9)\] or even \[360-90-5y-3-3y+9\] I am not sure what type of answer it wants.. I fi it needs to be simplified it would be \[276-8y\]

TheSmartOne (thesmartone):

If it needs**

TheSmartOne (thesmartone):

I am thinking it should be 1.2 hours...

TheSmartOne (thesmartone):

Yes if you need that in fraction that would be 204/170

TheSmartOne (thesmartone):

I am not sure...

OpenStudy (aum):

Let the distance between Donna's home and the mountains be d. Let 's' be the speed on her forward trip. The speed on her return trip is (s + 16) Speed = distance / time time = distance / speed Time for forward trip = d / s = 8 --- (1) Time for return trip = d / (s + 16) = 6 ---- (2) From (1): d = 8s. Put this in (2): 8s / (s + 16) = 6 8s = 6(s + 16) = 6s + 96 8s - 6s = 96 2s = 96 s = 48 d = 8 * s = 8 * 48 = 384 miles.

OpenStudy (amistre64):

looks like aum did all the work for you already ...